# Bundle-specific Tractogram Distribution Estimation Using Higher-order Streamline Differential Equation

### v1

@@ Abstract
Tractography traces the peak directions extracted from fiber orientation distribution (FOD) suffering from ambiguous spatial correspondences between diffusion directions and fiber geometry, which is prone to producing erroneous tracks while missing true positive connections. The peaks-based tractography methods “locally” reconstructed streamlines in ‘single to single’ manner, thus lacking of global information about the trend of the whole fiber bundle. In this work, we propose a novel tractography method based on a bundle-specific tractogram distribution function by using a higher-order streamline differential equation, which reconstructs the streamline bundles in ‘cluster to cluster’ manner. A unified framework for any higher-order streamline differential equation is presented to describe the fiber bundles with disjoint streamlines defined based on the diffusion tensor vector field. At the global level, the tractography process is simplified as the estimation of bundle-specific tractogram distribution (BTD) coefficients by minimizing the energy optimization model, and is used to characterize the relations between BTD and diffusion tensor vector under the prior guidance by introducing the tractogram bundle information to provide anatomic priors. Experiments are performed on simulated Hough, Sine, Circle data, ISMRM 2015 Tractography Challenge data, FiberCup data, and in vivo data from the Human Connectome Project (HCP) data for qualitative and quantitative evaluation. The results demonstrate that our approach can reconstruct the complex global fiber bundles directly. BTD reduces the error deviation and accumulation at the local level and shows better results in reconstructing long-range, twisting, and large fanning tracts.
{IEEEkeywords}
Diffusion MRI, Tractography, Bundle-specific tractography distribution, High-order streamline differential equation.
@@ 1 Introduction
[…]
TRACTOGRAPHY based on diffusion weighted magnetic resonance imaging (dMRI) is a potentially useful way of revealing the trajectories of white matter and structural connectome of the human brain in vivo [1, 2, 3, 4]. Typically, conventional tractography aims to integrate voxel-scale local fiber orientation extracted from fiber orientation distribution (FOD) to infer global connectivity, called FOD-based tractography [5], which faces challenges of producing large amounts of false-positives fibers and omitting true-positives fibers. These challenges have been widely discussed in schematic representations or theoretical arguments [6, 7, 8]. In general, it is usually attributed to the spatially ambiguous correspondences between diffusion directions and fiber geometry on voxel, local and global levels [1, 9, 10].
In recent decades, numerous efforts have been devoted to accurately estimating fiber trajectories, which are mainly reflected in diffusion modeling techniques and tractography strategies. For the diffusion modeling, high angular resolution diffusion imaging (HARDIs), such as constrained spherical deconvolution [11] and high-order tensor [12], are proposed to characterize multiple fibers within one voxel to overcome the limitation of DTI model. However, diffusion orientations are ambiguous when the asymmetric curvatures of the underlying fiber bundles are in the range of the voxel size [13, 3].
In recent years, asymmetric fiber orientation distributions (AFODs) have been proposed to address the problem of asymmetric fiber geometry. S. N. Sotiropoulos et al. [14] estimated fiber dispersion using Bingham distributions to represent continuous distributions of fiber orientations, centered on the main orientation, and captured anisotropic dispersion. Based on the geometric interpretation, Reisert et al. [15] derived a continuity condition that should be preserved for valid AFODs. Cetin et al. [13] proposed an asymmetric orientation distribution function (ODF) using a cone model in a voxel-wise manner. Considering the local surroundings of a voxel and using intervoxel information to derive potential fiber patterns is another approach used to estimating fiber geometry [16, 17, 18]. For instance, the estimation of AFODs based on a spherical deconvolution approach uses a set of symmetric and asymmetric basis functions by adding neighborhood continuity components [16]. However, these asymmetric models are good at representing such specific fiber geometry and simply plugging these AFODs into a typical tractography paradigm continues to face ambiguous spatial correspondences. Voxel-level or local-level diffusion models face challenges in resolving global fiber trajectory reconstruction with complex tractograms.
For tractography strategies, current tractography methods are based on the peaks extracted from FODs or AFODs, which we named peak-based tractography. Typical deterministic tractography tracks the peaks of the maximum diffusion direction, which inevitably accumulates error in the process. The probabilistic tractography algorithm obtains the fibers by randomly extracting directions from the FOD, leading to the production of substantial amounts of false-positive fibers [19, 20, 21, 22, 23]. Filter-based tractography [24, 25] can optimize the signal prediction error to iteratively reduce tractography biases when tracing the maximum diffusion direction. The signal is examined at each new position, and the filter recursively updates the underlying local model parameters to indicate the direction in which to propagate tractography. Filter-based methods cannot capture global fiber trajectory information to guide filtering and only optimize the fiber trajectory at the local level. Global optimization based tractography approaches [26, 27, 28, 29, 30, 31] reconstruct the fiber pathways, which aim to maximize the global energy of the vector field and the fiber structure solutions to prevent the fiber trajectory from being affected by local fiber direction errors. However, they can always find the maximal energy path between two seeds, and whether these optimal fibers exist is difficult to confirm. All these peak-based tractography algorithms reconstruct the bundles in a single streamline to single streamline manner from a seed point to an endpoint, which we referred to as ‘single to single’ tractography manner (Fig. 1a). From the view of the fiber bundle, we treat these ‘single to single’ tractography method as ‘local’ tractography methods. Substantial methodological innovation, such as directly reconstructing the fiber bundle from the starting region to the ending region, which we referred to as ‘cluster to cluster’ tractography manner, is a trend used to resolve these challenges. For example, Cottaar et al. [32] modeled fiber density and orientation using a divergence-free vector field to encourage an anatomically-justified streamline density distribution along the cortical white/gray-matter boundary while maintaining alignment with the diffusion MRI estimated fiber orientations. Aydogan et al. [33] proposed a novel propagation-based tracker that was capable of generating geometrically smooth curves using higher-order curves by using the more flexible parallel transport tractography (PTT) for curve parametrization. Based on the commonly accepted anatomical prior in which the fibers impossibly originate or terminate in the white matter and the hypothesis that fibers in the white matter show the form of non-intertwined streamlines, our previous works proposed a fiber trajectory distribution (FTD) function defined on the neighborhood voxels by using a ternary quadratic polynomial-based streamline differential equation [34]. The FTD can reveal continuous asymmetric fiber trajectory and showed an advantage over current methods. However, the FTD is still a local reconstruction of the fiber trajectory at the neighborhood voxel level.
@@
In this paper, we directly reconstruct a bundle-specific tractogram, i.e., a fiber bundle between two regions in a ‘cluster to cluster’ (Fig. 1b) instead of a voxelwise or local model combining the ‘single to single’ manner. To describe this global tractogram, we define a bundle-specific tractogram distribution (BTD) function based on any higher-order streamline differential equation in the measured diffusion tensor vector field. The optimization model is reconstructed with the measured diffusion orientations and BTD function. At the global level, the tractography process is to parameterize as BTD coefficients combining voxel location by minimizing the energy of optimization model. The relations between BTD and diffusion tensor vector are described under the prior guidance by introducing the tractography atlas as anatomic priors [35, 36]. This paper is organized as follows: The methods section defines the BTD based on higher-order streamline differential equation, which is resolved by global constraints in the diffusion tensor vector field. The experiments section presents the comparison results of the algorithms on three simulated datasets and the ISMRM 2015 Tractography Challenge dataset as well as the in vivo dataset of HCP [1, 37, 38]. The discussion section provides the conclusions of this work.
Figure 1: Tractography with different manners.
@@ 2 Methods
In its most basic form, a tractography algorithm takes two arbitrary points (voxels) of interest as input, labeled as a seed and a target, and yields the most likely trajectory on the given diffusion tensor vectorial field. Let $\Omega\subset{R^{3}}$ be the diffusion tensor vectorial field, and ${S_{U,V}}(t)\subset\Omega$ denotes a path that is parameterized by $t\in[0,T]$ connecting $U$ with $V$. Let $P(S)$ denote the probability of paths $S$ representing an anatomically genuine fiber trajectory in diffusion native space $\Omega\subset{R^{3}}$, which can be defined as,
@@
where ${V_{\max}}(s(t))$ is the principal diffusion direction of $s(t)$. In a sense, this is a “greedy” algorithm, as it tries to find the optimal fiber trajectory. The inherent unreliability and inaccuracy of deterministic fiber tracking have driven the introduction of new tracking algorithms. Stochastic tractography algorithms [41, 42] probabilistically generate the tracing directions based on the fiber conditional probability density function and the probability of direction within each voxel is commonly defined as ODF at the center of a voxel. Filter tractography algorithms use their current model state to obtain the predicted signal and combine the measured signal to propagate forward in most consistent direction. While shortest path methods aim to find an optimal path by computing the maximum energy between two seeds. In this paper, we are interested in the particular case of computing the diffusion tensor field on a Riemannian manifold [42, 43], which is a potential under the form $p\left({s(.),\dot{s}(.)}\right)=\sqrt{\dot{s}{{(.)}^{T}}{\rm M}(s(.))\dot{s}(.)}$ describing an infinitesimal distance along the fiber path $S$ relative to the metric tensor $M$ (symmetric definite positive). In this situation, finding the curve connecting two points that globally minimizes the energy [44] is a shortened path called a geodetic [45, 46]. In [47], a variational model is proposed based on Hamilton-Jacobi-Bellman, in which an infinite number of particles start from a given seed region evolving along the streamline orientation given by the gradient of the defined cost function to reach the seed target. In our recent work [17, 34], penalized geodesic tractography based on a Finsler metric with a global optimization framework is introduced to improve cortical connectomics. These methods are related to an optimal control problem and turn out to be very useful for establishing the connectivity of a single point on the seed region but provide little information about the connectivity between two regions. [48] attempted to introduce optimal mass transportation to describe the connectivity between two cerebral regions. Unfortunately, it is only a preliminary mathematical description without algorithmic implementation and validation. Herein, we will focus here on the bundle structure tractography connecting two cerebral regions in a ‘cluster to cluster’ manner (Fig. 1b). Considering two given regions ${\rho_{1}}$ and ${\rho_{2}}$, the optimal fiber bundle can be viewed as a superposition of non-intertwined fiber streamline cluster $S(t)$ starting from ${\rho_{1}}$ to ${\rho_{2}}$ along the diffusion tensor vector field, and the energy $P(S)$ assigned to the streamline cluster is defined as,
@@
where $S(0)={\rho_{1}}$, $S(T)={\rho_{2}}$ and $\dot{S}(t)$ is the diffusion vector field between ${\rho_{1}}$ and ${\rho_{2}}$. We first define the $S(t)$ and $\dot{S}(t)$, and $p\left({S(t),\dot{S}(t)}\right)$. In Section 2.1, we define the BTD function $S(t)$ on the diffusion vector field $\dot{S}(t)$ from ${\rho_{1}}$ to ${\rho_{2}}$ based on a higher-order streamline differential equation. For bundle continuity, we add spatial continuity constraint equation in Section 2.2. Then, the $S(t)$ solution can be simplified as estimation of BTD coefficients, detailed process is shown in Section 2.3.
@@ 2.1 Bundle-specific tractogram distribution function
Consider the diffusion field vector at position $(x,y,z)$ to be
@@
The tangent vector at point $(x,y,z)$ of fiber path ${s_{i}}$ equals the field vector $v(x,y,z)={\left[{{v_{x}}(x,y,z),{v_{y}}(x,y,z),{v_{z}}(x,y,z)}\right]^{T}}$, that is,
@@
The streamlines satisfy ${s_{i}}\cap{s_{j}}=0$, $i\neq j$, which is defined using the streamline differential equation,
@@
where ${a_{ijk}}$ is the coefficients of the polynomial. Combined with Eq. 7, the diffusion vector $v(x,y,z)$ is denoted,
$\begin{array}[]{l}v(x,y,z){\rm{=}}{\left[\begin{array}[]{l}\sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{x}{x^{i}}{y^{j}}{z^{k}}}}}\\ \sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{y}{x^{i}}{y^{j}}{z^{k}}}}}\\ \sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{z}{x^{i}}{y^{j}}{z^{k}}}}}\end{array}\right]^{T}}\\ {\rm{=}}A\cdot C(x,y,z)\end{array}$ (8)
@@
$\begin{array}[]{l}{C_{n}}(x,y,z)={({[{x^{n}},{x^{n-1}}y,{x^{n-1}}z,{x^{n-1}}{\rm{,}}\cdot\cdot\cdot,z{\rm{,}}1{\rm{}}]^{T}})_{\left({\sum\limits_{\Delta=0}^{n}{\frac{{(\Delta+1)*(\Delta+2)}}{2}}{\rm{,1}}}\right)}}\\ {\rm{=[}}{{\rm{F}}_{n}}(x,y,z){\rm{,}}{C_{n-1}}(x,y,z)]\\ \left\{\begin{array}[]{l}{C_{1}}(x,y,z)=[x,y,z,1],\\ {F_{n}}(x,y,z)=[{c_{m}},{\rm{}}\cdot\cdot\cdot{\rm{}},{c_{l}},\cdot\cdot\cdot{\rm{}}{c_{2}}{\rm{,}}{c_{1}}],{\rm{}}m=(n+1)(n+2)/2\\ {c_{l}}={x^{i}}{y^{j}}{z^{k}}{\rm{with}}\\ {\rm{}}l=i(n-\frac{{i-3}}{2})+j+1,{\rm{}}i=[0,1,\cdot\cdot\cdot,n],\\ j=[0,1,\cdot\cdot\cdot,n-i],{\rm{}}i+j+k=n\end{array}\right.\end{array}$ (9)
@@
$\begin{array}[]{l}{C_{2}}(x,y,z){\rm{=[}}{{\rm{F}}_{2}}(x,y,z){\rm{,}}{C_{1}}(x,y,z)]\\ {\rm{}}=[{x^{2}},{y^{2}},{z^{2}},x{y},x{z},yz,x,y,z,{1}]\\ \left\{\begin{array}[]{l}{C_{1}}(x,y,z)=[x,y,z,1]\\ {F_{2}}(x,y,z)=[{c_{6}},{c_{5}},{c_{4}},{c_{3}},{\rm{}}{c_{2}}{\rm{,}}{c_{1}}]{\rm{with}}\\ {\rm{}}{c_{6}}={x^{2}},{c_{5}}={y^{2}},{c_{4}}={z^{2}},{c_{3}}=x{y},{c_{12}}=x{z},{c_{1}}=y{z}\end{array}\right.\end{array}$

### v2

@@ Abstract
Streamline tractography locally traces peak directions extracted from fiber orientation distribution (FOD) functions, lacking global information about the trend of the whole fiber bundle. Therefore, it is prone to producing erroneous tracks while missing true positive connections. In this work, we propose a new bundle-specific tractography (BST) method based on a bundle-specific tractogram distribution (BTD) function, which directly reconstructs the fiber trajectory from the start region to the termination region by incorporating the global information in the fiber bundle mask. A unified framework for any higher-order streamline differential equation is presented to describe the fiber bundles with disjoint streamlines defined based on the diffusion vectorial field. At the global level, the tractography process is simplified as the estimation of BTD coefficients by minimizing the energy optimization model, and is used to characterize the relations between BTD and diffusion tensor vector under the prior guidance by introducing the tractogram bundle information to provide anatomic priors. Experiments are performed on simulated Hough, Sine, Circle data, ISMRM 2015 Tractography Challenge data, FiberCup data, and in vivo data from the Human Connectome Project (HCP) for qualitative and quantitative evaluation. Results demonstrate that our approach reconstructs complex fiber geometry more accurately. BTD reduces the error deviation and accumulation at the local level and shows better results in reconstructing long-range, twisting, and large fanning tracts.
keywords:
[…]
@@ 1 Introduction
Diffusion magnetic resonance imaging (dMRI) tractography (i.e., fiber tracking) is a useful tool for delineating white matter (WM) pathways and reconstructing structural connectome non-invasively in the in vivo human brain [1, 2, 3, 4, 5]. Typical tractography algorithms propagate streamlines through the vector field of fiber orientations extracted from fiber orientation distribution (FOD) functions estimated at each image voxel to reconstruct long-range connections [6]. This type of method is therefore often referred to as streamline tractography (ST, Fig. 1a) and is widely known to produce false-positive fibers and/or omit true-positive fibers [7, 8, 9], due to the ambiguous spatial correspondences between diffusion orientations and fiber geometry on the voxel-wise, local and global level [1, 10, 11]. Deterministic ST algorithms initiate each streamline from a seed voxel and then follow the peaks extracted from FODs to the termination region, which inevitably accumulates errors in the propagation process. Rather than using fixed fiber orientation(s), probabilistic ST algorithms randomly draw a direction based on the reconstructed FOD at each image voxel and perform the tracking many times using an iterative Monte Carlo approach, mapping all the possible pathways while producing a large number of false-positive fibers in the meanwhile [12, 13, 14, 15]. To reduce the generation of false-positive fibers, ST algorithms typically combine regions of interest (ROIs) selection strategies [16] or voxel-wise atlases [17] to analyze the generated fiber streamlines conform to anatomical regions.
Numerous signal modeling and tracking techniques have been proposed to improve the performance of streamline tractography. Early efforts focus on resolving multiple fibers within a voxel to solve the major problem of diffusion tensor imaging (DTI) which cannot resolve crossing and kissing fibers [18, 19, 20]. Nonetheless, these models still cannot accurately map fanning and bending fibers with asymmetric curvature in the range of the voxel size [3, 21]. Asymmetric FODs (AFODs) are consequently proposed to address this challenge [21, 22, 23]. A set of approaches leveraged local surrounding voxels and intervoxel correlation to characterize underlying fiber patterns  [24, 25, 26]. Unfortunately, AFODs only reduce the spatial ambiguity on the voxel and local level. Simply plugging them into a streamline tractography algorithm is not useful in mitigating the ambiguity on the global level for mapping complex tractograms. Additionally, many sophisticated tracking strategies have also been leveraged. For instance, “filtered tractography” reduces tracking biases by using an unscented Kalman filter to simultaneously fit the local the orientation distribution functions (ODF) model to diffusion signals and propagate a streamline in the most consistent direction  [27, 28]. “Parallel transport tractography” [29] generates geometrically smooth curves using the more flexible parallel transport frames and topographic regularity of connections. Nevertheless, the optimization only happens on the local level, without knowing any global information of the fiber trajectory.
Global tractography (GT) [30] methods aim to simultaneously reconstruct all tracks that can best describe the measured dMRI data (Fig. 1c). In this way, the estimation of local fiber orientations at each voxel is informed by surrounding voxels  [31, 32, 33]. For example, Jbabdi et al. [34] proposed a GT framework to guide probabilistic tracking, which can infer local directional parameters from large-scale connectivity in the presence of high local data uncertainty, based on a priori knowledge about regional correlations. GT offers improved agreement with the measured dMRI data as well as increased robustness to noise and imaging artifacts [35]. Even though GT algorithms map crossing fiber regions slightly more accurately, resultant bundle flow maps do not always represent the known fascicular topology [36]. The complexity of GT can be improved, at a cost of significantly increased computational burden, which is unacceptable or impossible in many applications.
Another category of tracking method is the recently proposed “bundle-specific tractography” (BST) [37]. Rather than mapping each streamline independently in ST or all streamlines simultaneously in GT, BST works on each fiber bundle of interest. Consequently, BST can incorporate anatomical and orientational prior knowledge of the tract into the tracking process (Fig. 1b). Specifically, BST first assigns fiber populations (e.g., peaks from the FOD) within a voxel to different fiber tracts, using a neural network  [38, 39] or an atlas, and performs classic streamline tractography only using the assigned fiber populations for each tract separately. BST results are more anatomically plausible, with more accurate delineation of intersected fiber tracts and larger spatial coverage to represent the full shape of bundles [40]. Nonetheless, current BST methods still use conventional deterministic or probabilistic tracking and suffer from pitfalls of streamline tractography.
Figure 1: Schematic representation of different tractography methods.
@@
To address this challenge, we propose to reconstruct all streamlines between the seeding and targeting region for a fiber tract of interest simultaneously, which is essentially semi-global or tract-wise global tractography. Specifically, we define a new bundle-specific tractogram distribution (BTD) function to describe the tract-wise tractogram based on any higher-order streamline differential equation in the measured diffusion tensor vector field. The optimization model is reconstructed with the measured diffusion orientations and BTD function. At the global level, the tractography process is to parameterize as BTD coefficients combining voxel location by minimizing the energy of the optimization model. The relations between BTD and diffusion tensor vector are described under the prior guidance by introducing the tractography atlas as anatomic priors.
Our proposed method is validated on three simulated datasets (Hough dataset, Sine dataset, and Circle dataset), the ISMRM 2015 Tractography Challenge dataset, FiberCup dataset, and the Human Connectome Project (HCP)  [41, 42] in vivo dataset. Results show that BTD can reconstruct complex fiber bundles, such as long-range, large twisting, and fanning tracts, and show better spatial consistency with fiber geometry. An open-source implementation of BTD is available at https://github.com/IPIS-XieLei/BTD-Tractography. Our contributions are summarized as follows:
  1. [(1)]
  2. i)
    A novel BTD function for fiber tractography to directly reconstruct the fiber trajectory is proposed.
  3. ii)
    A unified framework for any higher-order streamline differential equation is presented to describe the fiber bundles with disjoint streamlines defined based on the diffusion tensor vector field.
  4. iii)
    The fiber bundles are parameterized using the BTD coefficients, which are estimated by combining the priors and minimizing the energy on the diffusion tensor vector field.
  5. iv)
    Experimental results on Hough, Sine, Circle, FiberCup, ISMRM 2015 data, and HCP dataset show that the BTD is capable of reconstructing complex fiber bundles with long distances, large twists, and fan-shaped bundles, and shows better spatial consistency with the fiber geometry.
The rest of the paper is organized as follows. Section 2 defines the BTD based on a higher-order streamline differential equation. In Section 3, the experimental results on six datasets are presented. Section 4 discusses the results and future works, and Section  5 gives a summary of the proposed work.
@@ 2 Methods
In its most basic form, a tractography algorithm takes two arbitrary points (voxels) of interest as input, labeled as a seed and a target, and yields the most likely trajectory on the given diffusion vectorial field. Let $\mathcal{D}\subset{R^{3}}$ be the diffusion vectorial field, and ${S_{U,V}}(t)\subset\mathcal{D}$ denotes a path that is parameterized by $t\in[0,T]$ connecting $U$ with $V$. Let $\mathrm{P}(s)$ denote the probability of paths $s$ (Fig. 1a) representing an anatomically genuine fiber trajectory in diffusion native space $\mathcal{D}\subset{R^{3}}$, which can be defined as,
@@
where ${V_{\max}}(s(t))$ is the principal diffusion direction of $s(t)$. In a sense, this is a "greedy" algorithm, as it tries to find the optimal fiber trajectory. While shortest path methods aim to find an optimal path by computing the maximum energy between two seeds.
In this paper, we are interested in the particular case of computing the diffusion vectorial field on a Riemannian manifold [43], which is a potential under the form $p\left({s(.),\dot{s}(.)}\right)=\sqrt{\dot{s}{{(.)}^{T}}{\rm M}(s(.))\dot{s}(.)}$ describing an infinitesimal distance along the fiber path $S$ relative to the metric tensor $M$ (symmetric definite positive), which is usaually defined by the diffusion tensor [44]. In this situation, finding the curve connecting two points that globally minimizes the energy [45] is a shortened path called a geodetic [46, 47]. In [48], a variational model is proposed based on Hamilton-Jacobi-Bellman, in which an infinite number of particles start from a given seed region evolving along the streamline orientation given by the gradient of the defined cost function to reach the seed target. In our recent work [25, 49], penalized geodesic tractography based on a Finsler metric with a global optimization framework is introduced to improve cortical connectomics. These methods are related to an optimal control problem and turn out to be very useful for establishing the connectivity of a single point on the seed region but provide little information about the connectivity between two regions. Daducci et al. [50] attempted to introduce optimal mass transportation to describe the connectivity between two cerebral regions. Unfortunately, it is only a preliminary mathematical description without algorithmic implementation and validation.
Herein, we will focus here on the bundle structure tractography connecting two cerebral regions in a "cluster to cluster" manner. Considering two given regions ${\rho_{1}}$ and ${\rho_{2}}$, the optimal fiber bundle can be viewed as a superposition of non-intertwined fiber streamline cluster $S(t)$ starting from ${\rho_{1}}$ to ${\rho_{2}}$ along the diffusion vectorial field, and the energy $\mathrm{E}(S)$ assigned to the streamline cluster is defined as,
@@
where $S(0)={\rho_{1}}$, $S(T)={\rho_{2}}$ and $\dot{S}(t)$ is the diffusion vectorial field between ${\rho_{1}}$ and ${\rho_{2}}$. We first define the $S(t)$ and $\dot{S}(t)$, and $p\left({S(t),\dot{S}(t)}\right)$. In Section 2.1, we define the BTD function $S(t)$ on the diffusion vectorial field $\dot{S}(t)$ from ${\rho_{1}}$ to ${\rho_{2}}$ based on a higher-order streamline differential equation. For bundle continuity, we add spatial continuity constraint equation in Section 2.2. Then, the $S(t)$ solution can be simplified as estimation of BTD coefficients, detailed process is shown in Section 2.3.
@@ 2.1 Bundle-specific tractogram distribution function
Let the diffusion vectorial field at position $(x,y,z)$ be
@@
i) The tangent vector at each point $(x,y,z)$ of fiber path ${s_{i}}$ equals the field vector $v(x,y,z)$, that is,
@@
ii) The streamlines satisfy ${s_{i}}\cap{s_{j}}=0$, $i\neq j$, which is defined using the streamline differential equation,
@@
where ${a_{ijk}}$ is the coefficients of the polynomial. Combined with Eq. 4 and Eq. 7, the diffusion vector $v(x,y,z)$ is denoted,
$\begin{array}[]{l}v(x,y,z){\rm{=}}{\left[\begin{array}[]{l}\sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{x}{x^{i}}{y^{j}}{z^{k}}}}}\\ \sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{y}{x^{i}}{y^{j}}{z^{k}}}}}\\ \sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{a_{ijk}^{z}{x^{i}}{y^{j}}{z^{k}}}}}\end{array}\right]^{T}}\\ {\rm{=}}A\cdot C(x,y,z){\rm{}}\end{array}$ (8)
@@
$\begin{array}[]{l}{C_{n}}(x,y,z)={({[{x^{n}},{x^{n-1}}y,{x^{n-1}}z,{x^{n-1}}{\rm{,}}\cdot\cdot\cdot,z{\rm{,}}1{\rm{}}]^{T}})_{\left({\sum\limits_{\Delta=0}^{n}{\frac{{(\Delta+1)*(\Delta+2)}}{2}}{\rm{,1}}}\right)}}\\ {\rm{=[}}{{\rm{F}}_{n}}(x,y,z){\rm{,}}{C_{n-1}}(x,y,z)],\end{array}$ (9)
where ${C_{1}}(x,y,z)=[x,y,z,1]{\rm{}}$; ${F_{n}}(x,y,z)=[{c_{m}},\cdot\cdot\cdot{\rm{}},{c_{l}},\cdot\cdot\cdot{\rm{,}}{c_{1}}]$ with $m=(n+1)(n+2)/2$ and ${c_{l}}={x^{i}}{y^{j}}{z^{k}}$ where $l,i,j,k$ satisfy with $l=i(n-\frac{{i-3}}{2})+j+1$, $i=[0,1,\cdot\cdot\cdot,n]$, $j=[0,1,\cdot\cdot\cdot,n-i]$, and $i+j+k=n$.
@@
$\begin{array}[]{l}{C_{2}}(x,y,z){\rm{=[}}{{\rm{F}}_{2}}(x,y,z){\rm{,}}{C_{1}}(x,y,z)]\\ {\rm{}}=[{x^{2}},{y^{2}},{z^{2}},x{y},x{z},yz,x,y,z,{1}]\\ \left\{\begin{array}[]{l}{C_{1}}(x,y,z)=[x,y,z,1]{\rm{}}\\ {F_{2}}(x,y,z)=[{c_{6}},{c_{5}},{c_{4}},{c_{3}},{\rm{}}{c_{2}}{\rm{,}}{c_{1}}]{\rm{with}}\\ {\rm{}}{c_{6}}={x^{2}},{c_{5}}={y^{2}},{c_{4}}={z^{2}},{c_{3}}=x{y},{c_{12}}=x{z},{c_{1}}=y{z}\end{array}\right.\end{array}$